### v2

# Personalization Toolkit: Training Free Personalization of Large Vision Language Models

Abstract
Large Vision Language Models (LVLMs) have significant potential to provide personalized assistance by adapting to the unique needs and preferences of individual users. The personalization of LVLMs has emerged as a field that focuses on customizing models to recognize specific object instances and provide tailored responses. However, current methodologies depend on time-consuming test-time training for each user and object, which proves to be impractical. This paper introduces a novel, training-free approach to LVLM personalization by leveraging pre-trained vision foundation models to extract distinct features, retrieval-augmented generation (RAG) techniques to recognize instances in the visual input, and visual prompting methods. Our model-agnostic vision toolkit enables flexible and efficient personalization without the need for extensive retraining. We demonstrate state-of-the-art results, surpassing conventional training-based approaches, and set a new benchmark for LVLM personalization.
1 Introduction
Figure 1: Illustration of the personalization task and our PeKit. A reference image is introduced to the LVLM with information and possible context. The LVLM should later be able to answer questions about the introduced object using only the name of the object in the query. Our approach, PeKit, extracts patch-level features from the reference image and stores them in a memory module, $\mathcal{M}$. During personalized inference, our retrieval module, $\mathcal{R}$, queries $\mathcal{M}$ to detect the object. PeKit then informs the LVLM via a visual prompt, providing the name and possible context.
Large Vision Language Models (LVLMs) [23, 22, 7, 41, 20, 1, 36] have demonstrated impressive capabilities in reasoning about visual content and answering visual questions across various domains. This suggests a great potential for deployment as visual assistants that can aid users in their daily lives. However, current LVLMs are designed to provide generic, user-independent responses and recognize objects at the category level. This limits their ability to refer to specific instances by their unique names. Figure 1 shows an example of LVLM generic inference - i.e. giving a generic response to a user query.
The task of personalizing vision-language models was recently introduced by [2] to enable LVLMs to recognize specific object instances and answer relevant questions accordingly. However, existing approaches [2, 25] require training the model for each personalized object, which can be computationally expensive and limits the model’s ability to incrementally learn new personalized concepts.
In this work, we build upon the strengths of pre-trained vision foundation models, the emerging capabilities of LLMs in in-context learning [9] and retrieval-augmented generation (RAG) [17]. Our training-free approach localizes instances using open-world object detectors [24, 16, 26] and stores reference instance-level features in memory banks, alongside their name and context. During inference, our retrieval module queries the memory bank and visual-prompts the LVLM. By storing only feature embeddings, we avoid privacy concerns derived from training on user images. An overview of our approach is shown in Figure 1.
In addition to existing benchmark datasets, we evaluate our approach using a video personalization dataset, converted to an LVLM personalization format, and further augmented with a challenging evaluation set, which reflects a realistic personalization scenario and highlights the difficulty of the personalization task. It also shows the effectiveness of patch-level features for robust instance personalization. We also discuss the limitations of existing benchmark datasets and highlight opportunities for future research.
The key contributions of our approach are: 1) We demonstrate that LVLM personalization can be achieved without training, enabling fast and efficient deployment. 2) We propose a flexible approach that leverages pre-trained vision foundation models, RAG, and visual prompting to personalize any LVLM. 3) We achieve state-of-the-art results on three personalization benchmarks, outperforming existing approaches. 4) We create a challenging evaluation set that pushes the limits of existing approaches and highlights opportunities for future research.
In the following, we discuss the related work in Section 2 and present our vision toolkit for LVLMs personalization in Section 3, evaluate our approach in Section 3 and conclude in Section 5.
2 Related Work
Personalization of Large Language Models (LLMs). LLMs personalization [40] enhances user-specific interactions without extensive fine-tuning. Approaches like prompt tuning, in-context learning, and Retrieval-Augmented Generation (RAG) offer dynamic personalization without additional model storage [21, 4, 42]. Our method avoids storing new models or adapters and is training-free.
Text-to-Image Personalization. Personalizing text-to-image generation has been explored extensively, with early methods like Textual Inversion [10], Dreambooth [31], and HyperDreamVbooth [32] requiring training for each identity, leading to scalability issues. Recent approaches [33, 39, 13] avoid test-time fine-tuning by pretraining for personalization. However, we argue that, while personalized image generation does require training, LVLMs can be personalized without changing the large language model.
Large Vision Language Models Personalization The task of personalized vision language models was introduced in MyVLM [2] by training a concept head for specific objects on top of the CLIP cls token. In a similar fashion as Dreambooth [31], MyVLM employs rare tokens to encode personalized concepts. This may have unintended consequences for language assistants, and requires optimizing the LLM caption loss for personalized conversations. Yo’LLaVA [25] adds extra tokens to the LLM head for each personalized object, learning concept tokens to describe them. Although it surpasses MyVLM in performance, the additional tokens create a challenging incremental classification problem [8]. Both models need test-time training for each new concept, limiting them to personalizing one concept at a time. Two recent works [12, 27] attempt to eliminate test-time training. While Hao et al. [12] approaches the task by large-scale pretraining on personalized conversations, PLVLM [27] pretrains an alignment module to incorporate the CLS token and image embeddings from DINOv2 [26] into the LLM. Hao et al. [12] includes a reference image in the input to the LVLM, which may limit the capability to process multiple images and hinder in-context learning. PLVLM [27] relies on querying for a specific personalized object, limiting its applicability to only VQA. Our approach applies to both VQA and captioning (general and personalized) and is the first training-free personalization solution, featuring a modular, plug-and-play design that eliminates the need for retraining or adaptation.
Vision Foundation Models. Our approach uses vision foundation models, which have become powerful generalist models through large-scale pretraining [3]. Vision-language models like CLIP [29] and Align [15] map objects to higher-level semantics with contrastive pretraining. Self-supervised models like DINO [6, 26] detect variations within the same class. Grounding DINO [24] redefines object detection as phrase grounding, and SAM [16] excels in image segmentation.
Visual Prompting Visual prompting uses visual cues such as bounding boxes or arrows to guide Vision-Language Models. CLIP [28] interprets these marks to modify its cls token embedding accordingly [34]. Set of Mark Prompting [37] integrates GPT-4V with visual prompts using tools like MaskDINO [19], SAM [16], Semantic SAM [18], and SEEM [43]. ViPLLaVA [5] enhances LLaVA [23] to follow visual prompts by tuning on GPT-4V-labeled data. Contrastive Region Grounding (CRG) [35] improves LLaVA’s focus on objects by contrasting token probabilities with and without target object masking. Our experiments show that LLaVA and other LVLMs can describe objects accurately with proper instruction and context. Training-free methods like CRG [35] can further enhance attention if needed.
3 Approach
This section outlines our personalization toolkit, coined as PeKit, for enabling any LVLM to perform personalized object detection and answer generation. We employ a three-stage pipeline: View Extraction to extract robust object-level features from reference images and store them in a memory module, Personalized Objects Retrieval to identify objects in the query image, and Personalized Answer Generation via visual prompting. We refer to Fig 1 for an illustration of our approach.
3.1 Preliminary
We consider a set $P$ of all personalized objects introduced to a given LVLM. Each object $p\in P$ is associated with a set of reference images $\{I_{p}\}$, a name or identifier $n_{p}$ and optionally $c_{p}$ a context of the object. Our objective is to generate a personalized response for all images containing $p$ during inference, while producing a general caption for any other image that does not contain any of the personalized objects. The LVLM, e.g., LLaVA [23] typically takes as input an image $I_{p}$, a text query $Q$ and additional text as context or instruction.
3.2 Training-free View Extraction
Existing LVLM personalization techniques depend on image-level representations of the objects’ training views [2, 25], which can lead to overfitting to the background of each object in the reference images, particularly for training-based approaches. To avoid such a bias, our method localizes the object in the image and extracts only its corresponding features. We achieve this by utilizing an open-vocabulary segmentation network $F_{\text{ext}}$ to extract object-level masks $S_{p}$ based on each object’s generic category $k_{p}$ which can be deducted from the name or the context [^1] We refer to the Appendix for further discussion. .
${}S_{p}=F_{\text{ext}}(I_{p},k_{p}).$ (1)
We construct the average embedding vector $\mathbf{e}_{p}$ for object $p$ by average pooling of the embedding vectors produced by the image encoder $F_{\text{emb}}$ on the image $I_{p}$ over the region defined by the object-level mask $S_{p}$:
$\mathbf{e}_{p}=\operatorname{AvgPool}(F_{\text{emb}}{(I_{p}),S_{p})})\in\mathbb{R}^{D_{h}}.$ (2)
Considering $N$ reference images, we concatenate all object embedding vectors $\mathbf{e}_{p}^{i}$ (of the object $p$ pooled over the $i$-th reference image $I_{p}^{i}$) into a matrix $E_{p}=\left[\mathbf{e}_{p}^{1},\dots,\mathbf{e}_{p}^{N}\right]\in\mathbb{R}^{D_{h}\times N}$.
Memory module. After extracting the personalized objects’ embeddings, we store each object’s relevant properties in our memory module. The memory module is represented by a set $\mathcal{M}$ of object-specific entries:
=$\mathcal{M}=\{(E_{p},(n_{p},c_{p}))\}_{p\in P},$ (3)
where $n_{p}$ is the identifier or the name of the personalized object $p$, and $c_{p}$ is the context of the object, which can contain prior knowledge such as characteristics, background story or even relation to other personalized objects. When the number of personalized objects scales, the memory module $\mathcal{M}$ is easily converted into a Vector database, where nearest neighbor approximate search is deployed to retrieve instances matching a given query [11] ensuring the efficiency and scalability of our approach.
3.3 Personalized Object Retrieval
During inference, our goal is to determine whether a personalized object is present in the provided image $I$. We use any available object proposal technique $F_{\text{prop}}$ to generate a set of proposals (e.g., bounding boxes) $O=\{o_{i}\}_{i}=F_{\text{prop}}(I)$ for potential object occurrences within the image $I$. Then for each proposal $o_{i}$, we calculate its object-level average embedding vector:
$\mathbf{e}_{o_{i}}=\operatorname{AvgPool}(F_{\text{emb}}{(I),o_{i}}).$ (4)
We define the retrieval module $\mathcal{R}$ that takes an object embedding vector $\mathbf{e}_{o_{i}}$ and retrieves the memory entry $(E_{j},(n_{j},c_{j}))$ for a matching object $j$ as:
$\mathcal{R}(\mathbf{e}_{o_{i}})=\begin{cases}(E_{j},(n_{j},c_{j})),\\ \quad\text{if }j=\mathop{\mathrm{argmax}}_{l}\left\{\operatorname{sim}(E_{l},\mathbf{e}_{o_{i}})\right\}\\ \quad\text{and }\operatorname{sim}(E_{j},\mathbf{e}_{o_{i}})>\tau\\ \phi,\,\text{otherwise}\end{cases}$ (5)
where $\operatorname{sim}(E_{l},\mathbf{e}_{o_{i}})=\mathop{\mathrm{argmax}}_{k}\left\{\operatorname{sim}(\mathbf{e}_{l}^{k},\mathbf{e}_{o_{i}})\right\}$. We calculate the proposal’s similarity to the embeddings of the training views for all personalized objects. Any similarity measure (e.g., cosine similarity) can be employed for this purpose. We set a constant threshold $\tau$ to identify the personalized objects. We discard the object proposals in which no matching object is found by the retrieval module $\mathcal{R}$. Our method inherently supports the detection of multiple personalized objects (Fig 2).
3.4 Personalized Answer Generation
Once a personalized object is identified, our method generates captions specifically about that object, distinct from general captions a standard LVLM would produce. This involves emphasizing the detected object and incorporating prior knowledge about it. We achieve this through visual prompting by overlaying bounding boxes on the image and querying the LVLM to generate captions or answer questions focused on these objects. We use distinctive colors to differentiate recognized objects. We provide the LVLM with the object identifier $n_{j}$ (e.g., the instance name) and possibly a context $c_{j}$ for each personalized object. The LVLM incorporates this context and responds to queries using the given name $n_{j}$. For multiple personalized objects, we instruct the LVLM for each bounding box, name, and context. Figure 2 illustrates how the LVLM in our approach responds to the overlaid bounding boxes and our instruction. We refer to the Appendix for the exact prompt format. Note that generating answers using a context pool with varied information (e.g., monument history) is inherently supported by our method.
Figure 2: Multi-concept Personalization with Visual Prompting: PeKit handles multiple personalized objects without additional requirements. The LVLM in PeKit uses the specified bounding boxes (via visual prompts) and instance identities, with optional context (here imaginary), for personalized caption generation.
3.5 Choice of Vision Tools
We deploy GroundedSAM [30] as the open-vocabulary segmentation network $F_{\text{ext}}$ and ablate using GroundingDINO [24], where the mask is represented by the object’s bounding box. We use DINOv2 [26] as the image encoder $F_{\text{emb}}$ to extract patch-level features of the personalized objects. Image encoders trained with a self-supervised objective, such as DINOv2, produce distinctive features that improve the re-identification of personalized objects as we empirically show in the Appendix in a comparison with CLIP [28]. During inference, GroundingDINO [24] is queried with the generic term ‘object’ to generate object proposals.
3.6 Summary
To summarize our approach, there are two phases
Personalized object introduction: The user introduces an object to the LVLM with one or more reference images, name, and optional context. We extract robust patch-level features and store the object in our memory module $\mathcal{M}$ as $(E_{i},(n_{i},c_{i}))$.
Personalized Inference: Given an image and a question $Q$, we extract object proposals and query our retrieval module $\mathcal{R}$ with the bounding boxes features. If a match or multiple matches are found $\{(E_{i},(n_{i},c_{i}))\}$, we highlight the bounding box(es) on the query image with unique color(s) and instruct the LVLM to respond using $\{(n_{j},c_{j})\}$. If no match is found, the LVLM responds without further instructions.
4 Experiments
Figure 3: Our new evaluation set This-Is-My-Img based on This-Is-My dataset [38] : Training and validation sets examples for the category Reynard’s Work Chair.
This section evaluates our approach against existing methods for vision language personalization.
Datasets. MyVLM [2] dataset consists of 29 object categories and Yo’LLaVA [25] dataset includes 40 categories of objects, buildings, and people. It also features a VQA benchmark with multiple-choice questions (A/B). For both MyVLM and Yo’LLaVA datasets, we use only the reference images, discarding the ground-truth captions and negative images. This-Is-My-Img Dataset Yo’LLaVA and MyVLM datasets are simple and object-centric, lacking real-world complexity. To address this, we propose a new evaluation set for LVLM personalization based on the This-Is-My dataset [38], originally created for video-level detection of personalized instances reflecting real-world scenarios where objects may be partially visible or in the background. To make this dataset compatible with image-level training/evaluation, we process it as follows: We extract five reference frames per object from the training segments, ensuring the entire object is visible. For validation, we sample every 10th frame  [^2] Due to privacy concerns, some original video segments are unavailable, limiting our benchmark to 14 categories. See the Appendix for more details.. The validation frames are categorized into positives and negatives for each personalized object. Positives: frames containing the personalized object. Negatives: frames from the same validation segments where the object is not visible (denoted as Hard negatives) and frames from videos of other objects (Other). We further introduce an additional negative validation set (Fake) by generating images similar to the personalized objects frames with varied perspectives and environments using ChatGPT-4o (see Figure 3 for examples). We ensure that objects in the fake images differ in characteristics from the personalized objects. Our evaluation set challenges LVLM personalization methods with images containing multiple foreground objects, background elements, motion blur, and synthetic images [^3] This-Is-My-Img will be made publicly available.. We further extend This-Is-My-Img with a VQA evaluation set containing 70 multiple-choice (A/B) questions.
Implementation Details. Our method is generic regarding the choice of the LVLM model. We use LLaVA1.5-13B [22] as our primary LVLM model for consistency with previous works. We utilize Cosine Similarity with a constant threshold of $\tau$=0.75 for detecting personalized objects across all datasets (3.3). Refer to the Appendix for further details and results with automatic object-specific threshold using reference image similarity. We also report the compute and memory overhead of our personalization toolkit and compare them to the backbone in the Appendix.
Evaluation Metrics. We follow the metrics and notation of Yo’LLaVA [25] and MyVLM [2], reporting Positive Accuracy (Recall), Negative Accuracy (Specificity) and Weighted Accuracy for $n$ personalized objects. We further report Precision, defined as the ratio of true positive predictions to the total number of instances predicted as positive, averaged over $n$ personalized objects. We believe this is a more comprehensive evaluation recognizing the trade-off between the aforementioned metrics and precision. We report the accuracy (percentage of correct answers) for the VQA task. Following MyVLM, we report CLIPScore (measuring the similarity between the caption and the image) and personalization recall (assessing the LVLM’s ability to incorporate the personalized object’s name into the caption) on the MyVLM dataset.
MyVLM Dataset
Method/Metric Precision Accuracy
Positive Negative Weighted
MyVLM [2] - 96.6 90.9 93.8
Yo’ LLaVA [25] - 97.0 95.7 96.4
PeKit (G-DINO) 79.1 94.3 98.8 96.55
PeKit (G-SAM) 82.3 97.6 99 98.3
Yo’ LLaVA Dataset
Yo’ LLaVA [25] - 94.9 89.8 92.4
PeKit (G-DINO) 77 89.9 98.9 94.4
PeKit (G-SAM) 74.8 91 98.7 94.9
Table 1: Visual Recognition performance on MyVLM and YoLLaVA datasets. Our approach PeKit achieves state of the art performance and improves significantly over existing methods.
This-Is-My-Img Dataset
Method/Metric Precision Accuracy Avg.
Positive Negative
Other Hard Fake
MyVLM 8.0 88.1 14.6 4.7 54.2 33.9
Yo’LLaVA 42.1 87.1 84.4 61.9 61.4 67.3
PeKit 90.1 69.0 99.9 96.0 59.3 82.8
Table 2: Visual Recognition Performance (%) on This-Is-My-Img dataset. MyVLM and Yo’LLaVA tend to capture the dataset’s bias and produces many false positives when the images are taken from similar scenes regardless of the object’s presence.
4.1 Visual-Recognition
Table 1 presents the performance of our approach compared to current methods, MyVLM [2] and Yo’LLaVA [25], on their respective datasets. On the MyVLM dataset, our method PeKit achieves SOTA results, improving both positive and negative accuracy, with an average improvement of 1.9%. On the Yo’LLaVA dataset, our method significantly improves negative accuracy by 8.9% while having a lower positive accuracy than Yo’LLaVA [25], resulting in an average improvement of 2.5% on weighted accuracy, hence achieving best results on average. Our PeKit might miss a few challenging appearances of personalized objects but has a very low false positive rate, meaning we avoid incorrect recognition of personalized objects. We believe it is better to resort to generic captions than to provide incorrect personalized answers.
Method Metric
CLIPScore Personal. Recall
MyVLM Dataset
MyVLM [2] 27.6 94.76
PeKit (LLaVA) 30.2 97.1
Method Metric
VQA Accuracy
Yo’LLaVA Dataset
LLaVA [25] 89.9
Yo’LLaVA [25] 92.9
PeKit (LLaVA) 93.4
PeKit (InternVL) 95.9
This-Is-My-Img Dataset
LLaVA 72.8
Yo’LLaVA 67.1
PeKit (LLaVA) 77.1
PeKit (InternVL) 84.2
Table 3: Visual Question Answering Performance. PeKit achieves SOTA performance on the three datasets.
Next, we evaluate the performance on our introduced benchmark This-Is-My-Img in Table 2. MyVLM tends to provide positive responses to test images, resulting in very high positive accuracy but much lower negative accuracy ($14.6\%$ on other and $4.7\%$ on hard negatives). This indicates that MyVLM tends to capture the scene in which the object occurs rather than the specific object. Note that we followed exactly the training pipeline and requirements for positive and negative images of MyVLM. Yo’LLaVA has similar but less pronounced trends with high positive accuracy ($87.1\%$) and a moderate negative accuracy of $84.8\%$ on other negatives and $61.9\%$ on hard negatives. While PeKit achieves a lower positive accuracy, we demonstrate a more robust negative accuracy ($99.9\%$ and $96.0\%$ on all and hard negatives respectively). We achieve a precision of $90.1\%$ which is $48\%$ higher than Yo’LLaVA indicating that our approach refrains from providing wrong personalization. On the fake images, smaller differences appear between methods. The fake images sometimes contain very similar objects to the personalized object in terms of style, resulting in confusion by the methods, indicating room for future work. Yo’LLaVA and MyVLM detect the domain shift on some images more than PeKit. While PeKit is quite robust to different objects, it can be challenged on very similar fake and small (see limitations in Appendix). Overall, PeKit shows striking robustness and balanced behavior on both positive and negative accuracy achieving an average visual recognition performance of $82.86$ compared to 67.38 by Yo’LLaVA and 33.92 by MyVLM. Overall, this evaluation set illustrates the difficulty of the personalization task beyond object-centric images, moving closer to realistic scenarios for intelligent visual assistants. This opens the door for future research on this challenging benchmark.
4.2 Visual-Question Answering
In Table 3, we evaluate our method’s ability to answer questions about personalized objects using Yo’LLaVA [25] and This-Is-My-Img benchmarks. For completeness we also compare PeKit to MyVLM’s on the MyVLM dataset using CLIPScore [14] and Personalization Recall.
On the MyVLM dataset, PeKit achieves a higher performance than MyVLM method in generaing captions aligned with the images (CLIPScore) and incorporating the correct personalized concept names into its captions (Personalization Recall). On the Yo’LLaVA dataset our PeKit outperforms Yo’LLaVA without requiring training, special tokens, or modifications to the original LVLM.
Concerning This-Is-My-Img, all methods’ VQA performance experience a drop compared to the results on the Yo’LLaVA dataset, indicating the difficulty of our benchmark and room for future work. Due to this challenging nature and objects being sometimes in the background, we observe that Yo’LLaVA fails to improve over base LLaVA model. To the contrary, our PeKit achieves significant improvement of ($\sim 5\%$). This highlights our method strong ability in both visual recognition and question answering.
Figure 4: Qualitative Comparison to Yo’LLaVA: Yo’LLaVA’s prompt template requires specifying the personalized object’s identifier in the query (first row), limiting generalization since users must already know which objects are in the image. Using image-level embeddings can also cause confusion between similar objects (e.g., Alex vs. Alex’s bag). Adjusting the LLM’s head weights further harms captioning quality. PeKit achieves better captioning quality without any training. In each example, the “Query” shows Yo’LLaVA’s prompt with the concept identifier replaced by the object’s name and an added system prompt. PeKit uses a different template, detailed in the Appendix.
4.3 Ablations and Analysis
Choice of $F_{\text{ext}}$. Our method employs an open-vocabulary approach to localize personalized objects in reference images. In Table 1, we ablate the choice of an open-world object detector, G-DINO [24], compared to the open-world semantic segmentation model, G-SAM [30], for $F_{\text{ext}}$ in Eq. 1. The results show that our method outperforms previous methods in both cases. However, the more precise segmentation model, which extracts only patches of the object of interest, achieves the best accuracy on average. This improvement is more pronounced on the MyVLM dataset, where objects are less centered in the training views compared to the Yo’LLaVA dataset.
Other LVLMs. The modularity of our method allows for straightforward plug-and-play integration with any LVLM without tuning. We pair PeKit with InternVL2-26B [7] , a state-of-the-art LVLM. Table 3 shows VQA accuracy on the Yo’LLaVA and This-Is-My-Img datasets for PeKit with LLaVA [22] and InternVL [7]. With both backbones, our approach achieves SOTA results. The improved quality of InternVL [7] boosts VQA accuracy on the Yo’LLaVA dataset and achieves a significant $7\%$ boost on This-Is-My-Img.
Qualitative VQA Results Figure 4 shows examples of our PeKit compared to the Yo’LLaVA for VQA and personalized captioning tasks on This-is-my-Img dataset. As seen on the first row, Yo’LLaVA’s prompt template requires the query to include the target object’s concept identifier, making it unsuitable for general captioning tasks and tailored for Visual Question Answering. Besides, since Yo’LLaVA operates on image-level embeddings of reference views, it needs clutter-free object-centered reference views of the personalized objects. As seen on the second row, performance can decline if the personalized object is not in the foreground or if there are other objects/people interacting with the personalized object in the reference views. Besides, fine-tuning the last layer of the language model reduces the LVLM’s captioning capabilities for Yo’LLaVA.
In contrast, PeKit accurately identifies objects and uses visual prompting to guide LLaVA in answering the questions of ambiguous situations where multiple object are present in the image (e.g, Nikki’s car).
5 Discussion and Conclusion
In this work, we leverage vision foundation models combined with retrieval-augmented generation and visual prompting to build a training-free plug-and-play toolkit for the task of LVLM personalization called PeKit. We first show that our approach can surpass existing training-based methods while not requiring any tuning and not relying on any additional images other than those used to introduce the personalized concept. We further present a challenging benchmark that mimics more realistic scenarios and increases the complexity of the visual recognition challenge. While our approach shows robustness, there is room for improvement in future works. We believe that our PeKit sets a strong training-free baseline for the LVLM personalization task.